OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 3, -3, -9, ...

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

an = a1 + (n-1)d

OpenStudy (adilalvi):

-111

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

d = -6 in this case because your terms are being subtracted by 6 each time, so a(20) = 3 + (-6)(19)

OpenStudy (anonymous):

2.5 qt/min = ____ gal/h a. 60 c. 72 b. 37.5 d. 105

OpenStudy (adilalvi):

you welcome..

OpenStudy (anonymous):

? WHAT

OpenStudy (anonymous):

you didnt do anything lol

OpenStudy (adilalvi):

what you mean i didnt do nothing?

OpenStudy (anonymous):

all you said was -111 you didn't show any work how are we suppose to believe your answer

OpenStudy (adilalvi):

3,-3 add -6 to -3 and so on like -3+(-6)=-9 add -6 t0 -9.................

OpenStudy (adilalvi):

@ayoo_cuti123 ..?

OpenStudy (anonymous):

Oooo Lol Ok Sorry Geesh